# Emulating 2D Materials with Magnons

### v1

@@ I Introduction
The isolation of single layers of carbon in 2004 essentially gave birth to the experimental field of two-dimensional (2D) materials [novoselov_electric_2004], and immediately prompted further study of the quasi-relativistic nature of charge carriers in graphene [katsnelson_chiral_2006, li_observation_2007, guinea_energy_2010]. The catalog of 2D materials grew to include a zoo of electronic phases, including semiconductors [podzorov_high-mobility_2004] and insulators [novoselov_two-dimensional_2005] at first, but later more exotic interacting phases [ugeda_characterization_2016, cao_unconventional_2018, zeng_thermodynamic_2023]. Among the later species to join the zoo are 2D magnetic materials [han_graphene_2014, huang_layer-dependent_2017, gong_discovery_2017]. For example, ferromagnetism resembling both Ising-like and XY spins can be found within the family of chromium halides [liu_thickness-dependent_2019, bedoya-pinto_intrinsic_2021] and antiferromagnetism can be found in nickel halides [liu_vapor_2020, bikaljevic_noncollinear_2021]. However, much of this research goes beyond the equilibrium states of 2D magnets, and focuses also on novel magnetization dynamics.
@@
In magnetically ordered solids, including 2D materials, the low-energy spin excitations are usually wave-like collective excitations called spin waves (whose quanta are magnons – the terms will be used interchangably). Magnons are promising as charge-free information carriers for low-power computing devices in a field of study referred to as magnonics [kruglyak_magnonics_2010, serga_yig_2010, chumak_magnon_2015]. Unique applications of magnons include signal guiding using broken time-reversal symmetry and inference tasks that exploit their inherent nonlinearities [wang_inverse-design_2021, papp_nanoscale_2021, zenbaa_universal_2025, korber_pattern_2023]. Magnons in 2D and layered magnets bear similarities to their electronic counterparts and can host, for example, Dirac points [schneeloch_gapless_2022], topological gaps [chen_topological_2018], and flat bands [riberolles_chiral_2024]. Although some properties of these states are attractive, their applications seem distant because of their high frequency (on the order of THz) and limited number of experimentally readily accessible probes.
@@
One strategy of controlling magnons is given by magnonic crystals, analogous to the more well-known photonic crystals [krauss_two-dimensional_1996, joannopoulos_photonic_1997, nikitov_spin_2001]. By patterning a magnetic material into an artificial lattice, one may modify dispersion relations and open band gaps at the Brillouin zone (BZ) edges [chumak_current-controlled_2009, wang_nanostructured_2010, tacchi_band_2011]. This is a powerful control scheme for magnon transport that may find application in next-generation microwave and information technologies [chumak_magnonic_2017]. In this study, we outline a general strategy towards mapping between a few 2D materials and magnonic crystals supported by magnetization dynamics simulations and simple models. Understanding this mapping allows engineering of magnonic band structures using principles from 2D electronics. This brings some of the physics of 2D materials to more accessible frequency and length scales in a system that allows the addition of tailored defects and other spatial inhomogeneities. Therefore we provide a simple experimentally feasible platform that allows to probe directly the emerging physics of 2D materials, while also providing a design strategy for functional devices at microwave frequencies.
@@ II Spin Dynamics Simulations
Phenomenologically, classical magnetization dynamics are described by the Landau-Lifshitz-Gilbert [LLG] equation [landau_theory_1935, gilbert_phenomenological_2004]:
@@
where $\mathbf{M}$ is the magnetization with magnitude $M_{S}$, $\gamma$ is the gyromagnetic ratio, and $\mathbf{H}_{\text{eff}}$ is an effective magnetic field having contributions from externally applied fields, the exchange interaction, crystalline anisotropies, the demagnetizing field, etc. This first term gives rise to the well-known precessional motion of the magnetization $\mathbf{M}$. $\alpha$, the contribution of Gilbert, is a phenomenological damping parameter that weakens precession over time. A magnetic material may be thought of as being composed of many volume elements, known as macrospins, whose internal magnetization profiles are roughly uniform over a lenthscale given by the exchange length and who each individually follow this equation of motion. It is possible to perform simulations based on this principle, in which $\mathbf{H}_{\text{eff}}$ is calculated for every macrospin and time is evolved in steps. This approach to numerical simulations, referred to as ’micromagnetism,’ [brown_micromagnetics_1963] is rich in phenomena and is particularly helpful in studying magnetization switching, complex geometries and ground states, and nonlinear/soliton dynamics. Here, we use the GPU-accelerated micromagnetics program MuMax3 [vansteenkiste_design_2014] to study spin waves. [Fig. 1(a,b)].
@@ II.1 Geometry
Any periodic magnetic structure can be considered a magnonic crystal. However, geometries like arrays of disks face a disadvantage in being coupled only by the relatively weak magnetic dipole-dipole interaction. One approach is to embed magnetic components in a different magnetic matrix so the components may be loosely thought of as resonators coupled through spin waves in the matrix [puszkarski_magnonic_2003, krawczyk_magnonic_2013, centala_compact_2023, yang_flatbands_2025]. Another approach uses the so-called ”anti-dot” lattice, in which a continuous thin film is patterned with an array of holes [neusser_anisotropic_2010, ulrichs_magnonic_2010, bali_high-symmetry_2012, gros_phase_2021, wang_observation_2023]. The anti-dot lattice is different because it cannot generally be thought of as a clear set of coupled resonators. For instance, a lattice with vanishingly small holes should approach the behavior of a pristine film. Note that similar geometries have been explored in the artificial spin ice community, mostly for their interesting ground states [qi_direct_2008, gartside_realization_2018].
In general, a spin wave dispersion relation depends upon the magnetic ground state. Even for pristine thin films, the dispersion is anisotropic for an in-plane magnetization due to the dipole-dipole interaction [stancil_magnetostatic_2009]. To emulate the isotropic dispersion that is natural for electrons in 2D, we choose an out-of-plane (OOP) ground state by applying an external field $B_{\text{ext}}||\hat{z}$ sufficient to saturate the thin film [Fig. 1(c)]. Alternatively, one could also assume magnetic thin films that have perpendicular magnetic anisotropy. We choose material parameters of the prototypical magnonics material yttrium-iron garnet (YIG), which is widely used in magnonic devices due to its low dissipation. For these parameters, we choose anti-dot spacings of 100’s of nm to balance experimental feasibility with large frequency band widths. Because the simulation is performed using a rectangular array of pixel-like macrospins, it is tremendously convenient to work with a rectangular unit cell, so we define a rectangular supercell containing two primitive cells [see Fig. 1(d)]. As a result, the BZ is halved and contains double the number of bands [Fig. 1(e)]. Note that special points in the hexagonal BZ ($\mathbf{K}$ and $\mathbf{K^{\prime}}$, for instance) now fall inside the rectangular BZ, as shown in Fig. 1(f). These are technically new points, but we will refer to them as $\mathbf{K}$ and $\mathbf{K^{\prime}}$ for simplicity.
@@
For numerical simulations, a rectangular unit cell tens of micromagnetic cells in width is tiled tens of times; these two sizes determine real-space and inverse-space resolution, respectively. The simulation contains only one micromagnetic cell in the $\hat{z}-$direction – this is a good approximation for very thin films, in which the exchange interaction penalizes $\hat{z}-$direction texturing and standing waves (see supplemental Sec. LABEL:Supp-sec:S_Geometry). A film thickness of $15\text{ nm}$ is chosen to ensure accuracy. After the structure is relaxed to its OOP equilibrium state with periodic boundary conditions in the $x$ and $y$ directions, excitation can be applied to probe the magnonic band structure.
@@ II.2 Excitation and analysis
An excitation meant to probe a band structure should couple to a lot of wavevectors and frequencies – in other words, it should resemble a $\delta$-function. Therefore, within a small region, a radio frequency (rf) pulse is applied of the form $B_{x}(t)=B_{0}\frac{\text{sin}\left[2\pi f(t-t_{0})\right]}{2\pi f(t-t_{0})}$, whose spectrum contains all frequencies between f-f and +f+f. $B_{0}$ is chosen to be small ($\approx 0.1\text{ mT}$ or less) to make sure we stay in the linear regime. We time-evolve the system under the LLG equation and record the magnetization state $\mathbf{m}(t,x,y,z)$ as a unit vector. Precession about $\hat{z}$ is assumed, so the complex field:
@@
is a convenient representation that will be used frequently. This field can be interpreted as a wavefunction in some sense. Assuming circular precession and neglecting damping, eigenmodes (by definition) should keep their shape $|\Psi(\mathbf{r},t)|^{2}$ and time-evolve only by a phase: $\partial_{t}\Psi(\mathbf{r},t)=i\omega\Psi(\mathbf{r},t)$ for a specific angular frequency $\omega=2\pi f$. To give a relevant example, the linearized Landau-Lifshitz equation, including an exchange interaction of strength $D_{\text{ex}}$ and an externally applied field $H_{z}(\mathbf{r})$ can be rewritten (see appendix XI.1) in a Schrödinger-like form assuming a $\hat{z}-$polarized ground state[landau_magnetism_nodate, herring_theory_1951, lim_ferromagnetic_2021]:
@@
The time evolution of the magnetization after excitation is written in terms of $\Psi(\mathbf{r},t)$. Because it is a result of the excitation, we will refer to it as a response. Bloch’s theorem can be used to extract the magnonic band structure from the response: the signal is folded into $\Psi(t,x,y,i,j)$ where $(i,j)$ are unit cell indices and $(x,y)$ label positions within the unit cell. Then, the signal is Fourier transformed along the $t$, $i$, and $j$ axes, resulting in the complex amplitude $\Psi(f,x,y,k_{x},k_{y})$ associated with frequency $f$ and crystal momentum $(k_{x},k_{y})$. This approach is similar to Ref. feilhauer_unidirectional_2023.
@@ III Magnonic band structures
As expected, a continuous film [see Fig. 2(a)] has a magnonic band structure equivalent to the zone-folded representation of a free dispersion – the ”nearly-free-magnon.” Small holes [see Fig. 2(b)] significantly change the dispersion but do not open spectral gaps. However, large holes open spectral gaps and bring about some curious features [see Figs.2(c) and (d)]: near $1.5\text{ GHz}$, the band structure resembles that of graphene. Near $2.5\text{ GHz}$, the band structure seems to have a second set of Dirac points, and additionally has a few flat bands that meet dispersive bands at the $\mathbf{\Gamma}$ point. The nature of the transition between Fig. 2(b) and 2(c) is not obvious; see supplemental Sec. LABEL:Supp-sec:S_MoreBandStructures for an expanded figure.
@@
The isotropic flat bands of this system are particularly interesting. A truly flat band is one which has zero group velocity everywhere; its excitations are immobile. Fig. 3 illustrates this immobility. A rectangular region is continuously excited at the first flat band frequency. In the non-flat cases [see Figs. 3(a) and (b)], the power radiates away as usual. However, the flat band state cannot propagate and is instead excited to $\approx 1000\times$ the density of its counterparts [see Fig. 3(c)]. This is a rare feature, making this system potentially useful in nonlinear magnonics experiments (magnon Bose-Einstein condensation, for instance, in which large magnon densities must be achieved) [demokritov_boseeinstein_2006, tiberkevich_excitation_2019]. Outside magnonics, interactions in flat bands on the kagome lattice is a popular topic [wang_quantum_2023] – given the inherent nonlinearity of magnons, it is possible that this may provide a platform for an analog of interacting phases when this localized state is excited to large magnon number [wu_flat_2007].
Interestingly, the reason behind localization here can be understood from existing principles in 2D materials. Figure 3(d) is a zoomed-in version of Fig. 3(c), plotted using a phase-sensitive scheme. This reveals that the excited mode closely resembles the famous localized electronic wavefunction of the kagome lattice, whose localization property can be understood as the result of destructive interference between neighboring sites [rhim_singular_2021, chen_visualizing_2023, multer_imaging_2023]. This spin wave mode is also similar to localized spin excitations observed in the flat bands of layered kagome magnets, which can be understood by the same means [chisnell_topological_2015, riberolles_chiral_2024].
@@
To understand the real nature of the excitations, $\Psi(x,y)$ can be examined at a few $(f,k_{x},k_{y})$ points. This cell-periodic part of the response can be interpreted as a Bloch function. The magnon band structure is plotted along a few paths in Fig. 4(a), and some Bloch-like functions for $\mathbf{\Gamma}$-point modes are plotted in Fig. 4(b). One finding is that the lowest set of bands is indeed graphene-like: this figure demonstrates the Dirac point more clearly, but perhaps more incriminating is the fact that $\mathbf{\Gamma}$-point modes $1$ and $2$ in Fig. 4(b) are given by acoustic and optical modes of the honeycomb lattice. Actually, there exists a tiny gap at the Dirac point due to the dipole-dipole interaction, which is not easy to resolve (see supplemental Sec. LABEL:Supp-sec:S_DipoleGap). This is interesting but will be the primary subject of another study; here, we focus on a simple picture of the band structure.
@@ IV Tight-binding-like Analysis
Taking a step back, a simple Heisenberg spin chain under an externally applied magnetic field has the hamiltonian [stancil_quantum_2009]
@@
for an externally applied field $B_{0}||\hat{z}$ and ferromagnetic exchange parameter $\mathcal{J}$ which couples neighboring spins $\langle ij\rangle$. This can be rewritten using spin raising and lowering operators $S^{\pm}\equiv S_{x}\pm iS_{y}$. The Holstein-Primakoff (HP) transformation [holstein_field_1940] can then be used to rewrite this problem with magnon operators, assuming total spin $s$ parallel to $\hat{z}$ and small magnon number:
@@
This is the generic form of a tight-binding (TB) model for a 1D atomic chain, except that operators refer to magnons in the HP formalism. This suggests that TB models with appropriate parameters can generically describe exchange-only systems with ferromagnetic ground states. The interpretation is the same when these operators act on macrospins in nanostructures [iacocca_reconfigurable_2016]. With this TB picture in mind, it is not a large stretch to expect that nonuniform spin wave modes may also be represented by operators in this model, as long as the hopping parameters $t_{ij}$ and energies $\varepsilon_{i}$ are properly adjusted. For example, an array of exchange-coupled elements may be described by an effective TB model in which $a_{i}^{\dagger}$ creates a magnon in a nonuniform mode $i$, mainly confined to a single magnetic element [shindou_chiral_2013]. To be more precise, the nonuniform mode operators can be thought of as superpositions of magnon operators on many macrospins, with new hopping parameters that are simple to calculate from old parameters. In this picture, the hamiltonian becomes:
@@
Motivated by the observation that the Bloch functions [see Fig. 4(b)] are made of $s$ and $p$ ’orbitals,’ we choose this basis for a TB model. Using only nearest-neighbor hoppings, we implement this model using the PythTB package [coh_python_2022] and adjust parameters by hand. Fit parameters are listed in table 1 in units of frequency. Figure 5(a) is the same band structure appearing earlier, and Fig. 5(b) is its corresponding model, displayed in a manner similar to the simulations, using basis orbitals shown in Fig. 5(c) and (d). More precisely, we plot the spectrum of a supercell extended in the $\hat{y}-$direction with semitransparent coloring, and only plot the $k_{x}$ extent of the rectangular BZ.
@@
It is remarkable that a TB model works well in an anti-dot lattice. The orbitals emerge naturally from the structure without any careful engineering. The fact that a simple model works suggests that this is a general feature of the geometry that has nothing to do with spin waves. Indeed, simply solving the Schrödinger equation for a massive particle in a potential resembling this anti-dot lattice can yield a similar band structure (see supplemental Sec. LABEL:Supp-sec:S_Schrodinger). This phenomenon can be understood from Eq. 3. In fact, as a representation of the simplicity of excitations in this system, similar band structures have been discussed and observed in electronic, photon/polariton, acoustic, and cold-atom systems [milicevic_type-iii_2019, gao_visualization_2024, wu_p_xy-orbital_2008, polini_artificial_2013, mangussi_multi-orbital_2020, ding_experimental_2019].
Although the model appears to work, there are some inconsistencies. At the $\mathbf{\Gamma}$ point of the lowest band, the dispersion in the simulation is linear due to the dipole-dipole interaction [stancil_magnetostatic_2009], which is explicitly ignored by the TB model. Small mid-band gaps are also observable in the simulations, but this is only an artifact of the imperfect reconstruction of the hexagonal lattice using pixels, which breaks its $C_{6}$ symmetry down to $C_{2}$. Such gaps are not expected for real lattices. In addition, the model is less accurate at high frequencies. This is because the real wavelength of spin waves becomes small compared to the lattice features. For larger hole sizes (smaller lattice features), these high-frequency bands approach a more TB-like form (see supplemental Sec. LABEL:Supp-sec:S_MoreBandStructures). Other subtle inconsistencies exist, in part due to the TB model’s exclusion of the dipole-dipole interaction.

### v2

@@ I Introduction
The isolation of single layers of carbon in 2004 essentially gave birth to the experimental field of two-dimensional (2D) materials [65], and immediately prompted further study of the quasi-relativistic nature of charge carriers in graphene [40, 49, 26]. The catalog of 2D materials grew to include a zoo of electronic phases, including semiconductors [69] and insulators [66] at first, but later more exotic interacting phases [87, 6, 105]. Among the later species to join the zoo are 2D magnetic materials [28, 32, 24]. For example, ferromagnetism resembling both Ising-like and XY spins can be found within the family of chromium halides [55, 3] and antiferromagnetism can be found in nickel halides [53, 4]. However, much of this research goes beyond the equilibrium states of 2D magnets, and focuses also on novel magnetization dynamics.
@@
In magnetically ordered solids, including 2D materials, the low-energy spin excitations are usually wave-like collective excitations called spin waves (whose quanta are magnons – the terms will be used interchangably). Magnons are promising as charge-free information carriers for low-power computing devices in a field of study referred to as magnonics [44, 78, 13]. Unique applications of magnons include signal guiding using broken time-reversal symmetry and inference tasks that exploit their inherent nonlinearities [94, 67, 104, 41]. Magnons in 2D and layered magnets bear similarities to their electronic counterparts and can host, for example, Dirac points [76], topological gaps [9], and flat bands [74]. Although some properties of these states are attractive, their applications seem distant because of their high frequency (on the order of THz) and limited number of experimentally readily accessible probes.
@@
One strategy of controlling magnons is given by magnonic crystals, analogous to the more well-known photonic crystals [42, 36, 63]. By patterning a magnetic material into an artificial lattice, one may modify dispersion relations and open band gaps at the Brillouin zone (BZ) edges [11, 96, 85]. This is a powerful control scheme for magnon transport that may find application in next-generation microwave and information technologies [12]. In this study, we outline a general strategy towards mapping between a few 2D materials and magnonic crystals supported by magnetization dynamics simulations and simple models. Understanding this mapping allows engineering of magnonic band structures using principles from 2D electronics. This brings some of the physics of 2D materials to more accessible frequency and length scales in a system that allows the addition of tailored defects and other spatial inhomogeneities. Therefore we provide a simple experimentally feasible platform that allows to probe directly the emerging physics of 2D materials, while also providing a design strategy for functional devices at microwave frequencies.
@@ II Spin Dynamics Simulations
Phenomenologically, classical magnetization dynamics are described by the Landau-Lifshitz-Gilbert [LLG] equation [46, 22]:
@@
where $\mathbf{M}$ is the magnetization with magnitude $M_{S}$, $\gamma$ is the gyromagnetic ratio, and $\mathbf{H}_{\text{eff}}$ is an effective magnetic field having contributions from externally applied fields, the exchange interaction, crystalline anisotropies, the demagnetizing field, etc. This first term gives rise to the well-known precessional motion of the magnetization $\mathbf{M}$. $\alpha$, the contribution of Gilbert, is a phenomenological damping parameter that weakens precession over time. A magnetic material may be thought of as being composed of many volume elements, known as macrospins, whose internal magnetization profiles are roughly uniform and who each individually follow this equation of motion. It is possible to perform simulations based on this principle, in which $\mathbf{H}_{\text{eff}}$ is calculated for each macrospin and time is evolved in steps. This approach to numerical simulations, referred to as ’micromagnetism,’ [5] is rich in phenomena and is particularly helpful in studying magnetization switching, complex geometries and ground states, and nonlinear/soliton dynamics. Here, we use the GPU-accelerated micromagnetics program MuMax3 [89] to study spin waves. [Fig. 1(a,b)]. We choose realistic material parameters of the prototypical magnonics material yttrium-iron garnet (YIG), which is widely used in magnonic devices due to its low dissipation. See SM sec. LABEL:Supp-sec:S_Geometry for details [1].
@@ II.1 Geometry
Any periodic magnetic structure can be considered a magnonic crystal. However, geometries like arrays of disks face a disadvantage in being coupled only by the relatively weak magnetic dipole-dipole interaction. One approach is to embed magnetic components in a different magnetic matrix so the components may be loosely thought of as resonators coupled through spin waves in the matrix [71, 43, 7, 101]. Another approach uses the so-called ”anti-dot” lattice, in which a continuous thin film is patterned with an array of holes [62, 88, 2, 25, 90]. The anti-dot lattice is different because it cannot generally be thought of as a clear set of coupled resonators. For instance, a lattice with vanishingly small holes should approach the behavior of a pristine film. Note that similar geometries have been explored in the artificial spin ice community, mostly for their interesting ground states [72, 21].
In general, a spin wave dispersion relation depends upon the magnetic ground state. Even for pristine thin films, the dispersion is anisotropic for an in-plane magnetization due to the dipole-dipole interaction [82]. To emulate the isotropic dispersion that is natural for electrons in 2D, we choose an out-of-plane (OOP) ground state by applying an external field $B_{\text{ext}}||\hat{z}$ sufficient to saturate the thin film [Fig. 1(c)]. Alternatively, one could also assume magnetic thin films that have perpendicular magnetic anisotropy. We choose anti-dot spacings of 100’s of nm to balance experimental feasibility with large frequency band widths. Due to the rectangular shape of micromagnetic cells, it is tremendously convenient to work with a rectangular unit cell. We define a rectangular supercell containing two primitive cells [see Fig. 1(d)]. As a result, the BZ is halved and contains double the number of bands [Fig. 1(e)]. Note that special points in the hexagonal BZ ($\mathbf{K}$ and $\mathbf{K^{\prime}}$, for instance) now fall inside the rectangular BZ, as shown in Fig. 1(f). These are technically new points, but we will refer to them as $\mathbf{K}$ and $\mathbf{K^{\prime}}$ for simplicity.
@@
For numerical simulations, a rectangular unit cell tens of micromagnetic cells in width is tiled tens of times; these two sizes determine real-space and inverse-space resolution, respectively. The simulation contains only one micromagnetic cell in the $\hat{z}-$direction – this is a good approximation for very thin films, in which the exchange interaction penalizes $\hat{z}-$direction texturing and standing waves (see supplemental Sec. LABEL:Supp-sec:S_Geometry [1]). A film thickness of $15\text{ nm}$ is chosen to ensure accuracy. After the structure is relaxed to its OOP equilibrium state with periodic boundary conditions in the $x$ and $y$ directions, excitation can be applied to probe the magnonic band structure.
@@ II.2 Excitation and analysis
An excitation meant to probe a band structure should couple to many wavevectors and frequencies – in other words, it should resemble a $\delta(\mathbf{r},t)$-function. Therefore, within a small region, a radio frequency (rf) pulse is applied of the form $B_{x}(t)=B_{0}\frac{\text{sin}\left[2\pi f(t-t_{0})\right]}{2\pi f(t-t_{0})}$, whose spectrum contains all frequencies between f-f and +f+f. $B_{0}$ is chosen to be small ($\approx 0.1\text{ mT}$ or less) to make sure we stay in the linear regime. We time-evolve the system under the LLG equation and record the magnetization state $\mathbf{m}(t,x,y,z)$ as a unit vector. Precession about $\hat{z}$ is assumed, so the complex field:
@@
is a convenient representation that will be used frequently. This field can be interpreted as a wavefunction in some sense. Assuming circular precession and neglecting damping, eigenmodes (by definition) should keep their shape $|\Psi(\mathbf{r},t)|^{2}$ and time-evolve only by a phase: $\partial_{t}\Psi(\mathbf{r},t)=i\omega\Psi(\mathbf{r},t)$ for a specific angular frequency $\omega=2\pi f$. To give a relevant example, the linearized Landau-Lifshitz equation, including an exchange interaction of strength $D_{\text{ex}}$ and an externally applied field $H_{z}(\mathbf{r})$ can be rewritten (see appendix XI.1) in a Schrödinger-like form assuming a $\hat{z}-$polarized ground state[45, 29, 52]:
@@
The time evolution of the magnetization after excitation is written in terms of $\Psi(\mathbf{r},t)$. Because it is a result of the excitation, we will refer to it as a response. Bloch’s theorem can be used to extract the magnonic band structure from the response: the signal is folded into $\Psi(t,x,y,i,j)$ where $(i,j)$ are unit cell indices and $(x,y)$ label positions within the unit cell. Then, the signal is Fourier transformed along the $t$, $i$, and $j$ axes, resulting in the complex amplitude $\Psi(f,x,y,k_{x},k_{y})$ associated with frequency $f$ and crystal momentum $(k_{x},k_{y})$. This approach is similar to Ref. 18.
@@ III Magnonic band structures
As expected, a continuous film [see Fig. 2(a)] has a magnonic band structure equivalent to the zone-folded representation of a free dispersion – the ”nearly-free-magnon.” Small holes [see Fig. 2(b)] significantly change the dispersion but do not open spectral gaps. However, large holes open spectral gaps and bring about some curious features [see Figs.2(c) and (d)]: near $1.5\text{ GHz}$, the band structure resembles that of graphene. Near $2.5\text{ GHz}$, the band structure seems to have a second set of Dirac points, and additionally has a few flat bands that meet dispersive bands at the $\mathbf{\Gamma}$ point. The nature of the transition between Fig. 2(b) and 2(c) is not obvious; see supplemental Sec. LABEL:Supp-sec:S_MoreBandStructures [1] for an expanded figure.
@@
The isotropic flat bands of this system are particularly interesting. A truly flat band is one which has zero group velocity everywhere; its excitations are immobile. Fig. 3 illustrates this immobility. A rectangular region is continuously excited at the first flat band frequency. In the non-flat cases [see Figs. 3(a) and (b)], the power radiates away as usual. However, the flat band state cannot propagate and is instead excited to $\approx 1000\times$ the density of its counterparts [see Fig. 3(c)]. This is a rare feature, making this system potentially useful in nonlinear magnonics experiments (magnon Bose-Einstein condensation, for instance, in which large magnon densities must be achieved) [15, 86]. Outside magnonics, interactions in flat bands on the kagome lattice is a popular topic [95] – given the inherent nonlinearity of magnons, it is possible that this may provide a platform for an analog of interacting phases when this localized state is excited to large magnon number [98].
Interestingly, the reason behind localization here can be understood from existing principles in 2D materials. Figure 3(d) is a zoomed-in version of Fig. 3(c), plotted using a phase-sensitive scheme. This reveals that the excited mode closely resembles the famous localized electronic wavefunction of the kagome lattice, whose localization property can be understood as the result of destructive interference between neighboring sites [73, 8, 61]. This spin wave mode is also similar to localized spin excitations observed in the flat bands of layered kagome magnets, which can be understood by the same means [10, 74].
@@
To understand the real nature of the excitations, $\Psi(x,y)$ can be examined at a few $(f,k_{x},k_{y})$ points. This cell-periodic part of the response can be interpreted as a Bloch function. The magnon band structure is plotted along a few paths in Fig. 4(a), and some Bloch-like functions for $\mathbf{\Gamma}$-point modes are plotted in Fig. 4(b). One finding is that the lowest set of bands is indeed graphene-like: this figure demonstrates the Dirac point more clearly, but perhaps more incriminating is the fact that $\mathbf{\Gamma}$-point modes $1$ and $2$ in Fig. 4(b) are given by acoustic and optical modes of the honeycomb lattice. Actually, there exists a tiny gap at the Dirac point due to the dipole-dipole interaction, which is not easy to resolve (see supplemental Sec. LABEL:Supp-sec:S_DipoleGap [1]). This is interesting but will be the primary subject of another study; here, we focus on a simple picture of the band structure.
@@ IV Tight-binding-like Analysis
Taking a step back, a simple Heisenberg spin chain under an externally applied magnetic field has the hamiltonian [83]
@@
for an externally applied field $B_{0}||\hat{z}$ and ferromagnetic exchange parameter $\mathcal{J}$ which couples neighboring spins $\langle ij\rangle$. This can be rewritten using spin raising and lowering operators $S^{\pm}\equiv S_{x}\pm iS_{y}$. The Holstein-Primakoff (HP) transformation [31] can then be used to rewrite this problem with magnon operators, assuming total spin $s$ parallel to $\hat{z}$ and small magnon number:
@@
This is the generic form of a tight-binding (TB) model for a 1D atomic chain, except that operators refer to magnons in the HP formalism. This suggests that TB models with appropriate parameters can generically describe exchange-only systems with ferromagnetic ground states. The interpretation is the same when these operators act on macrospins in nanostructures [33]. With this TB picture in mind, it is not a large stretch to expect that nonuniform spin wave modes may also be represented by operators in this model, as long as the hopping parameters $t_{ij}$ and energies $\varepsilon_{i}$ are properly adjusted. For example, an array of exchange-coupled elements may be described by an effective TB model in which $a_{i}^{\dagger}$ creates a magnon in a nonuniform mode $i$, mainly confined to a single magnetic element [81]. To be more precise, the nonuniform mode operators can be thought of as superpositions of magnon operators on many macrospins, with new hopping parameters that are simple to calculate from old parameters. In this picture, the hamiltonian becomes:
@@
Motivated by the observation that the Bloch functions [see Fig. 4(b)] are made of $s$ and $p$ ’orbitals,’ we choose this basis for a TB model. Using only nearest-neighbor hoppings, we implement this model using the PythTB package [14] and adjust parameters by hand. Fit parameters are listed in table 1 in units of frequency. Figure 5(a) is the same band structure appearing earlier, and Fig. 5(b) is its corresponding model, displayed in a manner similar to the simulations, using basis orbitals shown in Fig. 5(c) and (d). More precisely, we plot the spectrum of a supercell extended in the $\hat{y}-$direction with semitransparent coloring, and only plot the $k_{x}$ extent of the rectangular BZ.
@@
It is remarkable that a TB model works well in an anti-dot lattice. The orbitals emerge naturally from the structure without any careful engineering. The fact that a simple model works suggests that this is a general feature of the geometry that has nothing to do with spin waves. Indeed, simply solving the Schrödinger equation for a massive particle in a potential resembling this anti-dot lattice can yield a similar band structure (see supplemental Sec. LABEL:Supp-sec:S_Schrodinger [1]). This phenomenon can be understood from Eq. 3. In fact, as a representation of the simplicity of excitations in this system, similar band structures have been discussed and observed in electronic, photon/polariton, acoustic, and cold-atom systems [60, 20, 99, 70, 57, 16].
Although the model appears to work, there are some inconsistencies. At the $\mathbf{\Gamma}$ point of the lowest band, the dispersion in the simulation is linear due to the dipole-dipole interaction [82], which is explicitly ignored by the TB model. Small mid-band gaps are also observable in the simulations, but this is only an artifact of the imperfect reconstruction of the hexagonal lattice using pixels, which breaks its $C_{6}$ symmetry down to $C_{2}$. Such gaps are not expected for real lattices. In addition, the model is less accurate at high frequencies. This is because the real wavelength of spin waves becomes small compared to the lattice features. For larger hole sizes (smaller lattice features), these high-frequency bands approach a more TB-like form (see supplemental Sec. LABEL:Supp-sec:S_MoreBandStructures [1]). Other subtle inconsistencies exist, in part due to the TB model’s exclusion of the dipole-dipole interaction.